OpenStudy (iwanttogotostanford):

calc:

OpenStudy (iwanttogotostanford):

@jim_thompson5910 please work me through the steps- if you can

jimthompson5910 (jim_thompson5910):

Let's first find f(-x) \[\Large f(x) = x^{-4} + \cos(-2x)\] \[\Large f(x) = \frac{1}{x^{4}} + \cos(2x)\] \[\Large f(-x) = \frac{1}{(-x)^{4}} + \cos(2(-x))\] \[\Large f(-x) = \frac{1}{x^4} + \cos(-2x)\] \[\Large f(-x) = x^{-4} + \cos(-2x)\] \[\Large f(-x) = f(x)\] So that shows f(x) is at least even Notes: anything to an even power is positive (eg: (-2)^4 = 2^4) Also, cos(-x) = cos(x)

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (iwanttogotostanford):

yes!

jimthompson5910 (jim_thompson5910):

Now let's find `-f(x)` \[\Large f(x) = x^{-4} + \cos(-2x)\] \[\Large -f(x) = -(x^{-4} + \cos(-2x))\] \[\Large -f(x) = -x^{-4} - \cos(-2x)\] Which is clearly different, so \[\Large f(-x) \ne -f(x)\] meaning that f(x) is NOT odd

jimthompson5910 (jim_thompson5910):

yeah it's even but not odd

jimthompson5910 (jim_thompson5910):

I think the only case where a function is both even and odd is when f(x) = 0 for all x

OpenStudy (iwanttogotostanford):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you recall how to find the slope of any line through 2 given points? ie, do you remember the slope formula?

OpenStudy (iwanttogotostanford):

y=mx+b @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Slope Formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\]

jimthompson5910 (jim_thompson5910):

you wrote out slope intercept form

OpenStudy (iwanttogotostanford):

oops yeah !@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you see what the y coordinates of the points are? if so, tell me what they are

jimthompson5910 (jim_thompson5910):

hint: it's the second coordinate of each point example: (5,7) has a y coordinate of 7

jimthompson5910 (jim_thompson5910):

Use the slope formula to get... \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{0-(-2)}{12-9}\] \[\Large m = \frac{0+2}{12-9}\] \[\Large m = \frac{2}{3}\] Then we use this slope and a point, say (12,0), to find the equation y = mx + b 0 = (2/3)(12) + b 0 = 8 + b b = -8 So the equation of this line in slope intercept form is \[\Large y = \frac{2}{3}x+8\]

OpenStudy (iwanttogotostanford):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

math is more my specialty

jimthompson5910 (jim_thompson5910):

sorry for some reason, I made a typo. Instead of \[\Large y = \frac{2}{3}x+8\] it should be \[\Large y = \frac{2}{3}x-8\]

jimthompson5910 (jim_thompson5910):

g(x) = f(x-2) + 4 First let's find f(x-2). Replace every copy of x with (x-2) f(x) = x^3 + 2 f(x-2) = (x-2)^3 + 2 f(x-2) = (x-2)*(x-2)^2 + 2 f(x-2) = (x-2)*(x^2-4x+4) + 2 f(x-2) = x*(x^2-4x+4)-2*(x^2-4x+4) + 2 f(x-2) = x^3-4x^2+4x-2x^2+8x-8 + 2 f(x-2) = x^3-4x^2-2x^2+4x+8x-8 + 2 f(x-2) = x^3-6x^2+12x-6 Now let's tack on a +4 at the end f(x-2)+4 = x^3-6x^2+12x-6+4 f(x-2)+4 = x^3-6x^2+12x-2 So, g(x) = x^3-6x^2+12x-2

jimthompson5910 (jim_thompson5910):

I would use a graphing tool like desmos https://www.desmos.com/calculator and locate the highest and lowest points

OpenStudy (iwanttogotostanford):

could you show me how and work through it with me?? @jim_thompson5910

OpenStudy (iwanttogotostanford):

can you example?

jimthompson5910 (jim_thompson5910):

what is the domain of y = cos(x) ? any ideas?

OpenStudy (iwanttogotostanford):

all real numbers?

jimthompson5910 (jim_thompson5910):

yes, how about the range?

OpenStudy (iwanttogotostanford):

-1≤y≤1?

jimthompson5910 (jim_thompson5910):

yep, now multiply all sides by 10 and you get what?

OpenStudy (iwanttogotostanford):

not sure... what you get>?

jimthompson5910 (jim_thompson5910):

\[\Large -1 \le y \le 1\] \[\Large 10(-1) \le 10*y \le 10*1\] \[\Large -10 \le 10y \le 10\] So the range of y = 10cos(x) is \(\Large -10 \le y \le 10\)

OpenStudy (iwanttogotostanford):

so then how would i find the maximum and minimum values ??

jimthompson5910 (jim_thompson5910):

the min and max are based on the range I just wrote out

OpenStudy (iwanttogotostanford):

could i ask you an econ question real quick that is fairly easy?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I can't guarantee anything

OpenStudy (iwanttogotostanford):

its fine- just give it your best shot please:-) I think oysters and stopwatches so far but i have to pick another too

jimthompson5910 (jim_thompson5910):

yep so far so good

jimthompson5910 (jim_thompson5910):

think of another item you would see on store shelves

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (iwanttogotostanford):

ok... and what about this one?? thanks for realizing that with me!

jimthompson5910 (jim_thompson5910):

you posted the same question twice

jimthompson5910 (jim_thompson5910):

resource and firm sound right

jimthompson5910 (jim_thompson5910):

resource and part of the circular flow maybe?

OpenStudy (iwanttogotostanford):

are you pretty sure on that? i think so too just want to be 100% sure...

jimthompson5910 (jim_thompson5910):

no not 100%

jimthompson5910 (jim_thompson5910):

which ones do you think?

OpenStudy (iwanttogotostanford):

choices B and D..?

jimthompson5910 (jim_thompson5910):

autos are part of the product market, not factor

OpenStudy (iwanttogotostanford):

oops i meant B and E

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

households offer labor because people that live in the houses offer to work for pay households offer capital in the form of investing (eg stocks) households also offer land. A landlord can rent out his/her land to a company to use

jimthompson5910 (jim_thompson5910):

in summary: households offer land, labor and capital

OpenStudy (iwanttogotostanford):

so you would say A and D then right???

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (iwanttogotostanford):

I'm thinking D or E... so last 2 cup for grabs there is only 1 right answer

jimthompson5910 (jim_thompson5910):

"monopsony" means "one buyer"

OpenStudy (iwanttogotostanford):

i know and I'm think either D or E .. thoughts?

jimthompson5910 (jim_thompson5910):

D and E are very similar, but congress is not a single entity like the navy is. Congress is composed of 2 houses with 535 members total, plus other staff

OpenStudy (iwanttogotostanford):

ohhh so just the Navy then would be right??

jimthompson5910 (jim_thompson5910):

besides, the term "Uniform" means "same" or "unified" in a sense. Which makes sense because the military all needs to be on the same page

OpenStudy (iwanttogotostanford):

oh ok- thanks!! I thought so but wasn't sure

OpenStudy (iwanttogotostanford):

im not so sure about this one....^

jimthompson5910 (jim_thompson5910):

that would be revenue revenue is the money flow of selling products and services (this is before taxes and expenses)

OpenStudy (iwanttogotostanford):

the answer i chose probably isnt even right

jimthompson5910 (jim_thompson5910):

probably A but idk about this one

OpenStudy (iwanttogotostanford):

me neither... I'm not so sure

jimthompson5910 (jim_thompson5910):

households, aka average everyday people anyways I have to get going so have a good night

OpenStudy (iwanttogotostanford):

wait one more!! its last one i promise!